# Zhang, Liu, and Zhao

### v4

@@ 4 Numerical Experiments
For the simulated data, we set the true model of $Y$ as $Y=\overline{Y}(X)+\widetilde{\epsilon}$, where $\overline{Y}(\cdot)$ represents the mean model, and $\widetilde{\epsilon}$ denotes a random variable with a standard normal distribution independent of $X$. Throughout all experiments with simulated data, we generate feature data uniformly from $[-1,1]^{p}$, and ensure the mean model satisfies $\min_{x\in[-1,1]^{p}}\overline{Y}(x)\geq 5$ which keeps all samples of $Y$ positive. Unless otherwise noted, both training and test datasets contain 1000 samples.
@@
Methods implemented in numerical experiments Abbreviation Explanation of method PADR Piecewise-affine decision rule method in this paper LDR Linear decision rule method from Ban and Rudin (2018) RKHS-DR Kernel optimizer prediction method from Bertsimas and Koduri (2021) kNN, kernel, cart, rf Prescriptive methods from Bertsimas and Kallus (2019) based on $k$-nearest neighbiors, kernels, regression trees, and random forests respectively sof Prescriptive method from Kallus and Mao (2023) based on stochastic optimization forests PO-PA, PO-L PO methods with piecewise-affine prediction model and linear prediction model respectively SIMOPT Simulated optimal decision using true conditional distribution
We refer to the PADR model with two max-affine components consisting of $K_{1}$ and $K_{2}$ pieces respectively as PADR$(K_{1},K_{2})$. The hyperparameters $K_{1}$ and $K_{2}$ are searched within the range [0, 4] for the Newsvendor problems in Section 4.1 and Section 4.2, and [0, 6] for the product placement problem in Section 4.3. Regarding the ESMM algorithm for learning the PADR, initial points are generated uniformly from $[-50,50]^{d(K_{1}+K_{2})(p+1)}$, and the sampling sequence $\{N_{\nu}\}_{\nu}$ follows $N_{\nu}=40\nu$.The parameter $\varepsilon$ in Algorithm 1 is set to be either constant or shrinking, the latter case of which is referred to as the shrinking-$\varepsilon$ strategy. Specifically, the shrinking strategy sets the parameter a constant value $\varepsilon_{0}$ for the first $T_{0}$ iterations and then a reduced value $\varepsilon_{1}(<\varepsilon_{0})$ for the remaining ones. The ESMM algorithm produces the iterate point obtaining the minimum ERM cost as the output after $T$ iterations. For each instance, the algorithm is executed for $r$ rounds with random initialization, and we report the best performance unless otherwise noted. We set $T=10$ and $r=10$ in Sections 4.1 and 4.2, and $T=20$ and $r=5$ in Section 4.3. The reported running time of PADR is the total time of all rounds.
We clarify settings for benchmark methods listed in Table 4. We use a Gaussian kernel for kernel methods. For RKHS-DR, we adjust the diagonal of the kernel by values no larger than $10^{-3}$ so as to resolve numerical issues without undermining performance. For PO methods, we use $\ell_{2}$ loss in the prediction stage and solve the regression problem of PO-PA using the ESMM algorithm. We set the uniform scale limit $\mu=50$ in Sections 4.1 and 4.2, and $\mu=100$ in Section 4.3. The hyperparameters of all benchmarks methods listed in Table 4 are selected via a standard validation process. The performance of LDR, PO-L, and kernel is dominated by PADR, PO-PA, and other prescriptive methods in most cases, and thus their results are omitted from the report if necessary to avoid the distraction.
@@
Hyperparameters of each method Methods Hyperparameters PADR number of pieces (PADR); sampling rate parameters, shrinking-$\varepsilon$ strategy, and proximal parameter (ESMM) LDR N/A RKHS-DR kernel bandwidth, regularization parameter, penalty parameter kNN, kernel, cart, rf number of nearest neighbors (kNN); kernel bandwidth, regularization parameter (kernel); depth, minimum samples to split, minimum samples per leaf (cart and rf); sof choice of apx-soln or apx-risk method in Kallus and Mao (2023), depth, minimum samples to split, minimum samples per leaf PO-PA number of pieces (prediction model); sampling rate parameters, shrinking-$\varepsilon$ strategy, and proximal parameter (ESMM solving $\ell_{2}$ regression) PO-L N/A SIMOPT N/A
All methods are implemented in Python 3.9.7 on a MacBook Pro with a macOS Monterey system and an Apple M1 Pro chip. All programs involved are built and solved by the optimizer Gurobi, version 9.5.0. We provide the code on GitHub: https://github.com/zhangyiy0223/PADR.
@@ 4.1 Newsvendor problem: simulated data
where $c_{b}$ and $c_{h}$ are the unit back-order and holding costs respectively, and $(\cdot)_{+}$ denotes $\max\{\cdot,0\}$. We set $(c_{b},c_{h})=(8,2)$ and consider the following max-affine demand mean model with $k=1$ unless otherwise specified:
$\overline{Y}(X)=k\max\{5X_{1}-10X_{2},-10X_{1}+5X_{2},15X_{1}\}+10.$ (17)
@@
Newsvendor problem: basic results and running time
Since the optimal DR in this case is a max-affine function with three pieces, PADR$(K_{1},K_{2})$ with $K_{1}\geq 3$ has no approximation error in principle. Figure 1d shows that the costs of PADR with $(K_{1},K_{2})=(3,0)$ converge to SIMOPT as the training sample size $n$ increases from 50 to 1000. RKHS-DR converges slower than PADR; additionally, its performance with $n=1000$ is worse than that of PADR with $n=400$. Prescriptive methods, including kNN, cart, rf, and sof, are dominated by the PADR and RKHS-DR methods. For PO-PA, we observe that the prediction model provided by the PO-PA method is quite close to the true model $\overline{Y}(\cdot)$, so its substantial gap to SIMOPT could be mainly due to inconsistency between the prediction loss and Newsvendor cost with imbalanced $c_{b}$ and $c_{h}$. Numerical results with the parameter ratio $c_{b}:c_{h}$ varying from 1:9 to 9:1 exhibit similar patterns and thus are omitted. Figure 1e shows the training time of each method. The training time of RKHS-DR grows roughly exponentially when $n\geq 200$, whereas the time for PA-related methods remains stable and significantly shorter when $n=1000$, which is mainly credited to the enhanced and sampling techniques used in the ESMM algorithm.
@@
Performance of ESMM under various settings In (a), solid and dotted lines represent constant-$\varepsilon$ strategies; dashed lines represent shrinking-$\varepsilon$ strategies with $T_{0}=3$ and $\varepsilon_{1}=0$. In (b), experiments for $p=20$ are conducted on datasets using the sparse demand model introduced in Section 4.1.2. The y-axis is set on a log scale.
In Figure 1g, we provide results of the ESMM algorithm under different parameter settings that illustrate the effect of the enhanced sampling strategies. Figure 1f displays curves of sorted training cost of PADR provided by the ESMM algorithm under the constant-$\varepsilon$ and shrinking-$\varepsilon$ strategies over 30 rounds. Under the constant-$\varepsilon$ strategy, the algorithm converges to stationary solutions which can be strengthened to global optimality with $\varepsilon$ being sufficiently large. However, Figure 1f shows that the constant-$\varepsilon$ strategy typically yields poor performance over all 30 rounds with 10 iterations. We conjecture that this is because when the sample size is large, the algorithm requires engaging with the random selection among a huge number of active indices and thus hardly makes a valid update at each iteration. In contrast, the shrinking-$\varepsilon$ strategy enables flexible adjustments of active indices for early iterations with a large positive value of $\varepsilon$ and produces iterate points serving as a warm start; afterward, it accelerates the convergence with a decreased value of $\varepsilon$. Figure 1f shows that the shrinking-$\varepsilon$ strategy has significantly better performance which attains the global optimum among roughly half of the total rounds.
@@
(h) Sparse demand model
(i) Dense demand model
@@
Figure 1i displays the results. In the sparse case, forest-based prescriptive methods are robust to redundant features, as they only focus on informative features. When the dimension is less than 20, PADR method is competitive with comparable performance to sof and much less training time. It would be beneficial to employ an effective feature selection technique or regularization in the PADR method to improve the performance in the sparse case with a large feature dimension. In the dense case, PADR performs generally well, but its performance slightly deteriorates as $p$ increases due to growing estimation errors and optimization difficulties. We also observe that PADR with fewer pieces performs more stably in both cases despite the presence of approximation errors.
@@
We study the effect of nonlinearity of the demand with respect to the feature variables by varying the coefficient $k$ in the demand model (17) and report results in Figure 1k. It indicates that the performance of PADR and PO-PA is irrelevant to the nonlinearity level of the demand model due to the strong approximation ability to the nonlinear structure. In contrast, the performance of RKHS-DR and prescriptive methods deteriorates significantly as the nonlinearity level increases. These methods require significantly more samples to achieve the similar performance when the level of nonlinearity increases.
@@
(j) $p=2\phantom{xxxxxxx}$
(k) $p=20$
@@ 4.2 Newsvendor problem with nonconvex cost: real-world data
In this section, we conduct experiments on a real-world dataset, which consists of order records for a specific model of mobile phone from a Chinese electronic product supplier. The order record, spanning from September 01, 2021 to November 31, 2023, contains order quantities and time stamps from retailers. In order to create a production plan, the supplier needs to determine the packing quantity in each day to fulfill the order requests.
We first organize the individual order records into daily order quantities. The aggregated daily order quantities have been rescaled for confidentiality purposes and also normalized for pretreatment. The boxplots of daily orders in Figure 1n show that order quantities are larger on weekdays and during the second quarter of each month, which aligns with retailers’ ordering pattern.
@@
(l) Weekdays
(m) Month days
(n) Month quarters
@@
We set observations from the first year as the training set and the remaining data as the test set. Here we consider three types of features, namely weekday, month quarter, and order quantities in past days. The first two types of features are converted into dummy variables (7 for weekdays and 4 for month quarters) and set as basic features. For the third type of features, we incorporate order quantities in past days of different lengths as different instances in conducting the experiments. To determine the packing quantity in each day to fulfill the order requests, we formulate a Newsvendor problem with features listed above and report the numerical results in Table 4.2. It shows that PADR and forest-based prescriptive methods work well. Overall, the PADR method performs slightly better in most instances with features of different length.
@@
Test costs for newsvendor on a real-world dataset Features $p$ DR methods Prescriptive methods PO methods PADR RKHS-DR kNN cart rf sof PO-PA basic features 11 55.95 56.10 56.06 56.00 56.00 56.00 79.02 basic features + order quantities in past 1 day 13 42.67 51.53 50.25 61.72 49.54 48.70 68.18 basic features + order quantities in past 3 days 15 41.65 55.59 48.25 54.64 41.74 42.19 59.73 basic features + order quantities in past 5 days 17 41.40 65.17 49.59 56.74 40.90 41.83 52.39 basic features + order quantities in past 7 days 19 40.93 70.60 49.02 65.62 43.62 43.51 54.52 basic features + order quantities in past 14 days 26 41.03 78.37 46.81 58.39 42.45 42.48 52.10 basic features + order quantities in past 21 days 33 42.43 80.24 46.61 59.45 41.59 41.33 53.44 The lowest three costs in each row are highlighted in bold and color, with the lowest one being underlined and colored red.
We next consider a practical setting where the supplier incurs an additional cost associated with packing quantities in actual production operations, in order to balance the daily inventory cost and the packing cost. The cost is approximated by the following nonconvex cost function:
$C(z)=\min\{z,0.6z+0.8,0.4z+15.6\},$ (18)
Therefore, we consider a Newsvendor problem with an additional concave cost $C(z)$ in the objective. Since the concave cost introduces nonconvexity into the problem. we use the deterministic MM algorithm to solve the ERM problem for learning RKHS-based decision rule, referred to as RKHS-DR(MM). For prescriptive methods, we solve one-dimensional nonconvex problems by line search. We do not implement sof from Kallus and Mao (2023) in this nonconvex Newsvendor problem, as it requires a new development of second-order perturbation analysis of stochastic optimization, which is beyond the scope of this paper.
@@
Features $p$ DR methods Prescriptive methods PO methods
PADR RKHS-DR(MM) kNN cart rf PO-PA
basic features 11 74.28 81.89 84.52 81.56 81.56 90.48
basic features + order quantities in past 1 day 13 64.75 72.11 77.86 86.58 73.78 77.89
basic features + order quantities in past 3 days 15 64.10 68.29 75.58 78.31 64.99 73.61
basic features + order quantities in past 5 days 17 63.20 64.72 76.81 79.54 65.74 71.56
basic features + order quantities in past 7 days 19 62.01 65.35 75.38 85.95 68.35 69.97
basic features + order quantities in past 14 days 26 64.86 65.15 73.81 75.52 64.71 70.80
basic features + order quantities in past 21 days 33 65.65 65.15 73.81 75.52 65.73 71.02
@@
[…]

### v5

@@ 4 Numerical Experiments
For the simulated data, we set the true model of $Y$ as $Y=\overline{Y}(X)+\widetilde{\epsilon}$, where $\overline{Y}(\cdot)$ represents the mean model, and $\widetilde{\epsilon}$ denotes a random variable with a standard normal distribution independent of $X$. Throughout all experiments with simulated data, we generate feature data uniformly from $1,1]^{p}$, and ensure the mean model satisfies $\min_{x\in[-1,1]^{p}}\overline{Y}(x)\geq 5$ which keeps all samples of $Y$ positive. Unless otherwise noted, both training and test datasets contain 1000 samples.
@@
Methods implemented in numerical experiments Abbreviation Explanation of method PADR Piecewise-affine decision rule method in this paper RKHS-DR Kernel optimizer prediction method from Bertsimas and Koduri (2021) NN-DR Neural network-based decision rule method with two or three fully connected hidden layers LDR Linear decision rule method from Ban and Rudin (2018) kNN, KERNEL, CART, RF Prescriptive methods from Bertsimas and Kallus (2019) based on $k$-nearest neighbiors, kernels, regression trees, and random forests respectively SOF Prescriptive method from Kallus and Mao (2023) based on stochastic optimization forests PO-PA, PO-L PO methods with piecewise-affine prediction model and linear prediction model respectively SIMOPT Simulated optimal decision using true conditional distribution
@@
We clarify settings for benchmark methods listed in Table 4. For NN-DR, we use ReLU function ($\max\{\cdot,0\}$) as activation, and the number of nodes in each hidden layer is set to half the size of the preceding layer. In each experiment, the model is trained for 100 epochs over 3 independent runs using the Adam optimizer (Kingma and Ba 2014) to minimize the empirical cost with $\ell_{1}$-regularization. We use a Gaussian kernel for kernel methods. For RKHS-DR, we adjust the diagonal of the kernel by values no larger than $10^{-3}$ so as to resolve numerical issues without undermining performance. For PO methods, we use $\ell_{2}$ loss in the prediction stage and solve the regression problem of PO-PA using the ESMM algorithm. We set the uniform scale limit $\mu=50$ in Sections 4.1 and 4.2, and $\mu=100$ in Section 4.3. The hyperparameters of all benchmarks methods listed in Table 4 are selected via a standard validation process. The performance of LDR, PO-L, and KERNEL is dominated by PADR, PO-PA, and other prescriptive methods in most cases, and thus their results are omitted from the report if necessary to avoid the distraction.
[…]
Hyperparameters of each method Methods Hyperparameters PADR number of pieces (PADR); sampling rate parameters, shrinking-$\varepsilon$ strategy, and proximal parameter (ESMM) RKHS-DR kernel bandwidth, regularization parameter, penalty parameter NN-DR first-hidden-layer size, regularization parameter, learning rate, weight decay, batch size LDR N/A kNN, KERNEL, CART, RF number of nearest neighbors (kNN); kernel bandwidth, regularization parameter (KERNEL); depth, minimum samples to split, minimum samples per leaf (CART and RF); SOF choice of apx-soln or apx-risk method in Kallus and Mao (2023), depth, minimum samples to split, minimum samples per leaf PO-PA number of pieces (prediction model); sampling rate parameters, shrinking-$\varepsilon$ strategy, and proximal parameter (ESMM solving $\ell_{2}$ regression) PO-L N/A SIMOPT N/A
All methods are implemented in Python 3.9.7 on a MacBook Pro with a macOS Monterey system and an Apple M1 Pro chip. NN-DR is implemented in Pytorch version 2.2.2. All programs involved are built and solved by the optimizer Gurobi, version 9.5.0. We provide the code on GitHub: https://github.com/zhangyiy0223/PADR.
@@ 4.1 Newsvendor problem: simulated data
where $c_{b}$ and $c_{h}$ are the unit back-order and holding costs respectively, and $(\cdot)_{+}$ denotes $\max\{\cdot,0\}$. We set $(c_{b},c_{h})=(8,2)$ and consider the following max-affine demand mean model with $\omega=1$ unless otherwise specified:
$\overline{Y}(X)=\omega\max\{5X_{1}-10X_{2},-10X_{1}+5X_{2},15X_{1}\}+10.$ (17)
@@
Newsvendor problem: basic results and running time PO-L, KERNEL, and LDR are excluded in (a) for readability, whose costs are around 14, 11, and 9 respectively.
[…]
@@
(h) Time across $K_{1}$ and $K_{2}$
Performance of ESMM under various settings In (a), solid and dotted lines represent constant-$\varepsilon$ strategies; dashed lines represent shrinking-$\varepsilon$ strategies with $T_{0}=3$ and $\varepsilon_{1}=0$. In (b), experiments for $p=20$ are conducted on datasets using the sparse demand model introduced in Section 4.1.2. The y-axis is set on a log scale. In (c), each point represents the running time of PADR with $K_{1}$ and $K_{2}$ varying from 0 to 40 for the Newsvendor problem with $p=2$ and $n=1000$.
In Figure 1h, we provide results of the ESMM algorithm under different parameter settings that illustrate the effect of the enhanced sampling strategies. Figure 1f displays curves of sorted training cost of PADR provided by the ESMM algorithm under the constant-$\varepsilon$ and shrinking-$\varepsilon$ strategies over 30 rounds. Under the constant-$\varepsilon$ strategy, the algorithm converges to stationary solutions which can be strengthened to global optimality with $\varepsilon$ being sufficiently large. However, Figure 1f shows that the constant-$\varepsilon$ strategy typically yields poor performance over all 30 rounds with 10 iterations. We conjecture that this is because when the sample size is large, the algorithm requires engaging with the random selection among a huge number of active indices and thus hardly makes a valid update at each iteration. In contrast, the shrinking-$\varepsilon$ strategy enables flexible adjustments of active indices for early iterations with a large positive value of $\varepsilon$ and produces iterate points serving as a warm start; afterward, it accelerates the convergence with a decreased value of $\varepsilon$. Figure 1f shows that the shrinking-$\varepsilon$ strategy has significantly better performance which attains the global optimum among roughly half of the total rounds.
@@
Figure 1h illustrates the running time of PADR for multiple settings of $(K_{1},K_{2})$ with each value ranging from 1 to 40. The running time for 10 rounds remains within 100 seconds when $\max\{K_{1},K_{2}\}\leq 40$. An interesting observation is that, with the same total number of pieces (i.e., total number of parameters), the time consumption of PADR is relatively higher when its two components have more balanced piece numbers, and considerably lower when the decision rule consists of only one component (i.e., $K_{1}=0$ or $K_{2}=0$). This can be explained by the complex structure of the surrogate function for the PADR with balanced piece numbers.
@@
(i) Sparse demand model
(j) Dense demand model
@@
[…]
(k) Sparse demand model
(l) Dense demand model
Running time in Newsvendor problem with varying feature dimensions
Figure 1j and Figure 1l report the test cost and running time, respectively. In the sparse case, forest-based prescriptive methods are robust to redundant features, as they only focus on informative features. It would be beneficial to employ an effective feature selection technique in the PADR method to improve the performance in the sparse case with a large feature dimension. In the dense case, PADR performs generally well, but its performance slightly deteriorates as $p$ increases due to growing estimation errors and optimization difficulties. We also observe that PADR with fewer pieces performs more stably in both cases despite the presence of approximation errors. When $p<20$, PADR consistently achieves slightly better performance than benchmarks, while requiring substantially less training time. For instance, at $p=8$, PADR attains a 10% lower test cost compared to NN-DR and SOF, with a 90% less runtime than these methods. This highlights the efficiency of PADR in learning decision rules in settings with moderate feature dimensions.
@@
We study the effect of nonlinearity of the demand with respect to the feature variables by varying the coefficient $\omega$ in the demand model (17) and report results in Figure 1o. It indicates that the performance of PADR and PO-PA is irrelevant to the nonlinearity level of the demand model due to the strong approximation ability to the nonlinear structure. In contrast, the performance of RKHS-DR and prescriptive methods deteriorates significantly as the nonlinearity level increases. These methods require significantly more samples to achieve similar performance when the level of nonlinearity increases. NN-DR shows the robustness towards non-linearity for the demand model with $p=20$; however, additional experiments in Figure 1o reporting relative cost gaps to SIMOPT for PADR, RKHS-DR, and NN-DR with multiple dimension settings, show that the performance of RKHS-DR deteriorates as $p$ increases, while NN-DR becomes less sensitive to nonlinearity as $p$ increases. We speculate that when $p$ is small, the instability of gradient-based optimization becomes more pronounced, and this effect is further amplified under highly nonlinear demand functions.
To further test the effect of nonlinearity, we consider demand models with quadratic, cubic, and sine terms (in addition to a shared linear component). Results in Online Appendix 12 show consistent patterns across methods, validating PADR’s robustness to model misspecification under varying nonlinear structures.
@@
(m) $p=2\phantom{xxxxxxx}$
(n) $p=20$
(o) Relative Cost Gap ($\omega=3$)
@@ 4.2 Newsvendor problem with nonconvex cost: real-world data
In this section, we conduct experiments on a real-world dataset, which consists of order records for a specific model of mobile phone from a Chinese electronic product supplier. The order record, spanning from September 01, 2021 to November 30, 2023, contains order quantities and time stamps from retailers. In order to create a production plan, the supplier needs to determine the packing quantity in each day to fulfill the order requests.
We first organize the individual order records into daily order quantities. The aggregated daily order quantities have been rescaled for confidentiality purposes and also normalized for pretreatment. The boxplots of daily orders in Figure 1r show that order quantities are larger on weekdays and during the second quarter of each month, which aligns with retailers’ ordering pattern.
@@
(p) Weekdays
(q) Month days
(r) Month quarters
@@
We set observations from the first year as the training set and the remaining data as the test set. Here we consider three types of features, namely weekday, month quarter, and order quantities in past days. The first two types of features are converted into dummy variables (7 for weekdays and 4 for month quarters) and set as basic features. For the third type of features, we incorporate order quantities in past days of different lengths as different instances in conducting the experiments. Note that incorporating retailer’s past order quantities into the feature vector introduces temporal dependence, violating the i.i.d. assumption in Assumption (A1). This issue can be addressed using generalization bounds for non-stationary $\beta$-mixing processes, as established by Kuznetsov and Mohri (2017). Since only a fixed-length history is used in the feature, the resulting data sequence is $\beta$-mixing with finite memory, i.e., there exists a lag $a_{0}$ such that $\beta(a)=0$ for all $a>a_{0}$. Under such settings, Theorem 1 and Corollary 1 from Kuznetsov and Mohri (2017) ensure that the generalization bound still holds, and thus the excess risk analysis remains valid.
To determine the packing quantity in each day to fulfill the order requests, we formulate a Newsvendor problem with features listed above and report the numerical results in Table 4.2. The results show that PADR and NN-DR exhibit competitive performance. PADR outperforms NN-DR when $p\leq 15$, whereas NN-DR demonstrates greater robustness as the feature dimension increases.
@@
Test costs for Newsvendor on a real-world dataset Added features $p$ DR methods Prescriptive methods PO methods (all include basic features) PADR RKHS-DR NN-DR kNN CART RF SOF PO-PA None 11 55.95 56.10 56.74 56.06 56.00 56.00 56.00 79.02 + order quantities in past 1 day 13 42.67 51.53 45.13 50.25 61.72 49.54 48.70 68.18 + order quantities in past 3 days 15 41.65 55.59 42.46 48.25 54.64 41.74 42.19 59.73 + order quantities in past 5 days 17 41.40 65.17 41.78 49.59 56.74 40.90 41.83 52.39 + order quantities in past 7 days 19 40.93 70.60 40.27 49.02 65.62 43.62 43.51 54.52 + order quantities in past 14 days 26 41.03 78.37 41.04 46.81 58.39 42.45 42.48 52.10 + order quantities in past 21 days 33 42.43 80.24 40.12 46.61 59.45 41.59 41.33 53.44 The lowest three costs in each row are highlighted in bold and color, with the lowest one being underlined and colored red.
We next consider a practical setting in which the supplier incurs an additional cost due to packaging requirements in actual production operations, aiming to balance daily inventory and packing costs. The packing cost follows an incremental discount scheme (Tersine and Barman 1991), where the unit price decreases for quantities exceeding certain thresholds. To align with the rescaled order quantities, we specify the following piecewise affine concave cost function:
$C(z)=\min\{z,0.6z+8,0.4z+15.6\}.$ (18)
Accordingly, we consider a Newsvendor problem with an additional term $C(z)$ in the objective function. The inclusion of this concave term introduces nonconvexity into the optimization problem. we use the deterministic MM algorithm to solve the ERM problem for learning RKHS-based decision rule, referred to as RKHS-DR(MM). For prescriptive methods, we solve one-dimensional nonconvex problems by line search. As SOF from Kallus and Mao (2023) requires a new development of second-order perturbation analysis of stochastic optimization to accommodate the nonconvex term, we instead replace $C(z)$ by its affine piece $0.6z+8$ when computing weights for training samples. We refer to this linearized version as SOF(L).
@@
Added features $p$ DR methods Prescriptive methods PO methods
(all include basic features) PADR RKHS-DR NN-DR kNN CART RF SOF(L) PO-PA
None 11 74.28 81.89 80.67 84.52 81.56 81.56 82.50 90.48
+ order quantities in past 1 day 13 64.75 72.11 68.19 77.86 86.58 73.78 74.45 77.89
+ order quantities in past 3 days 15 64.10 68.29 66.62 75.58 78.31 64.99 66.54 73.61
+ order quantities in past 5 days 17 63.20 64.72 64.07 76.81 79.54 65.74 67.66 71.56
+ order quantities in past 7 days 19 62.01 65.35 62.25 75.38 85.95 68.35 69.23 69.97
+ order quantities in past 14 days 26 64.86 65.15 63.45 73.81 75.52 64.71 65.58 70.80
+ order quantities in past 21 days 33 65.65 65.15 63.33 73.81 75.52 65.73 66.01 71.02
@@
Table 4.2 demonstrates that PADR consistently outperforms other methods when $p\leq 17$, with performance gains exceeding 5% for $p\leq 13$, and achieves the lowest test cost at $p=19$. These findings highlight PADR’s effectiveness in addressing nonconvex contextual SP problems. On the other hand, NN-DR maintains superior performance when $p\geq 26$. Forest-based methods, however, perform relatively poorly in this nonconvex setting. The performance gap between SOF(L) and RF becomes more pronounced than in the convex case, since linearizing the concave cost introduces considerable model misspecification. Moreover, the persistently weak performance of PO-PA across both convex and nonconvex scenarios indicates that its sequential prediction and decision-making process may not be suitable for this real-world dataset.